OpenStudy (anonymous):

The directrix of y =1/8 x^2 is: A. x = -2 B. y = -2 C. y = 2

OpenStudy (amoodarya):

rewrite it as standard form

OpenStudy (amoodarya):

\[x^2=4py\\\]

OpenStudy (anonymous):

That how they wrote it

OpenStudy (amoodarya):

\[x^2=16y\\4p=16\\p=4\]

OpenStudy (amoodarya):

that was an example ! and what is yours ?

OpenStudy (amoodarya):

\[\frac{1}{8}x^2=y\\* 8\\x^2=8y\\x^2=4py\\\]

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

is the answer y=2